i from q_01 Cross-sectiona Surface > It is unnecessary to include this document in your homework submittal. Given: 1. The chaHints: 1. Exponential functions are ubiquitous in Electrical Engineering as well as all areas of engineering. You might needA = -8nC

B = 11nC

tau = 6usec

t_1 = 30usec
Answer #1

a (a q)A+ Be て- Gusec つ8- =V B llnC t 30usec 8nCt(11c)e5o)yc no =(-8+11e M 0stt ie 0sts30usec OveY l(-8+% 8+110)x6 3x5 ()fo t(f t-+usec 5S4 nC (4)Ss4nC - for t-usec t) =8+1e IS.9 (2) =6SlnC M)for t- 1Ssec )=(-8+1 )xl = X o -709nC 1(15)70nC vi t- 18usix)fx t-24s, 1)= (-8+112{»)xlo =-7.8nC 9(24)-7,8nC xlo 8+1le ( rt- sec, q= E&+ -7.87nC -7. 87n C (K)or t- 30uset=(-8+1e5) xldl-8n+lne n) dt 6)i)=dof) nx/ X dt 1lme )=11emA o1 for t- 3nse e xl 3) mA (for t-6use -Lle 0 67m A 0.6m A =-0.409 m4 JE) ) for t1use 6 -0409 mA ivfr t- 12uec i(t= -le{63j = -O osSimA 6 RI)-0055MA Vii) fot- 24psec,iE}= LIEG -0. 033mA 0r033 m A (ix) tor t7Asec,t -1 e 0.02 m A xlo -0.02 m A have tovethFeel free to ask if you have any query.Thank you.